# EMD: Explicit Motion Modeling for High-Quality Street Gaussian Splatting

### v1

@@ Abstract
Photorealistic reconstruction of street scenes is essential for developing real-world simulators in autonomous driving. While recent methods based on 3D/4D Gaussian Splatting (GS) have demonstrated promising results, they still encounter challenges in complex street scenes due to the unpredictable motion of dynamic objects. Current methods typically decompose street scenes into static and dynamic objects, learning the Gaussians in either a supervised manner (e.g., w/ 3D bounding-box) or a self-supervised manner (e.g., w/o 3D bounding-box). However, these approaches do not effectively model the motions of dynamic objects (e.g., the motion speed of pedestrians is clearly different from that of vehicles), resulting in suboptimal scene decomposition. To address this, we propose Explicit Motion Decomposition (EMD), which models the motions of dynamic objects by introducing learnable motion embeddings to the Gaussians, enhancing the decomposition in street scenes. The proposed EMD is a plug-and-play approach applicable to various baseline methods. We also propose tailored training strategies to apply EMD to both supervised and self-supervised baselines. Through comprehensive experimentation, we illustrate the effectiveness of our approach with various established baselines. The code will be released at: https://qingpowuwu.github.io/emdgaussian.github.io/.
@@
{strip}[Uncaptioned image]
Figure 1: Previous street Gaussian splatting methods find it challenging to accurately model the motion of dynamic objects, which leads to blurry reconstructions. With the introduction of the proposed Explicit Motion Decomposition (EMD), which improves the decomposition of dynamic object motion, the current method achieves significantly better reconstruction quality.
@@ 1 Introduction
Novel view synthesis for dynamic scenes is essential in autonomous driving, enabling a variety of applications such as simulation, testing, and validation of perception systems. Traditional simulators like CARLA (Dosovitskiy et al. 2017) and AirSim (Shah et al. 2018), while providing controlled environments, suffer from limited realism and require substantial manual effort in creating virtual scenes. Recent advances in neural rendering, particularly Neural Radiance Fields (NeRF) (Mildenhall et al. 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023), have emerged as promising alternatives for photorealistic scene reconstruction. These pioneering approaches excel in capturing complex geometries and appearances through implicit or explicit neural representations, and have been extended to dynamic scenes by incorporating an additional time dimension into their representations and learning deformation networks to model non-rigid motions (Park et al. 2021a; Pumarola et al. 2020; Li et al. 2021; Tretschk et al. 2020; Park et al. 2021b; Xian et al. 2020; Xu, Alldieck, and Sminchisescu 2021; Wang et al. 2022; Cleac’h et al. 2022; Cao and Johnson 2023; Rabich, Stotko, and Klein 2023). Despite advancements, reconstructing autonomous driving scenes remains difficult due to complex multi-object dynamics, expansive environments, and varied motion patterns.
@@
To tackle this challenge, existing works usually build upon dynamic NeRF and 3DGS frameworks, further separating autonomous driving scenes into static and dynamic components through two primary paradigms: Supervised methods utilize pre-trained models to acquire auxiliary conditions, such as segmentation masks from SAM (Kirillov et al. 2023), depth maps from DepthAnything (Yang et al. 2024a), and both depth and optical flow from Dynamo (Sun and Hariharan 2023), or 3D bounding boxes from various datasets. Representative works like StreetGaussian (Yan et al. 2024) demonstrate the effectiveness of using supervision signals for trajectory optimization and appearance modeling. In contrast, self-supervised methods achieve static-dynamic separation without explicit supervision, as shown by S3Gaussian (Huang et al. 2024), which leverages inherent motion cues. While both paradigms have shown promising results, their binary classification of scene elements as either static or dynamic overlooks the continuous spectrum of motion inherent in real-world street scenes.
@@
To better address the different motion patterns in street scenes, we propose an Explicit Motion Decomposition (EMD) module that can be easily integrated into existing supervised and self-supervised frameworks (Fig. 2). EMD improves scene decomposition by incorporating motion-aware feature encoding and dual-scale deformation modeling. Specifically, we enhance each Gaussian primitive with learnable motion embeddings to capture its motion characteristics and design a hierarchical deformation framework that separately manages fast, global motions and slow, local deformations. This design allows for more efficient analysis of complex street scenes with varying motion speeds.
To showcase the versatility of our approach, we conduct extensive experiments on the Waymo-NOTR dataset by integrating EMD with representative methods from both paradigms: StreetGaussian for supervised and S3Gaussian for self-supervised settings. Our main contributions include:
@@
  • We propose EMD, a new plug-and-play module that effectively addresses varying motion speeds in street scenes through explicit motion modeling.
  • We validate our method in both supervised and self-supervised settings, demonstrating consistent performance improvements across various evaluation protocols.
  • We carried out a thorough evaluation of the Waymo-NOTR dataset, showcasing substantial improvements across various settings, achieving better reconstruction quality in both full scenes (+1.81 PSNR) and vehicle-specific regions (+2.81 PSNR) compared to respective baselines.
@@ 2 Related Work
Neural Radiance Fields (NeRF) (Mildenhall et al. 2020) revolutionized novel view synthesis by introducing volumetric scene representation using MLPs(Taud and Mas 2018), and has been significantly improved through various extensions (Barron et al. 2021, 2022; Müller et al. 2022; Liu et al. 2020). However, these methods face challenges in dynamic scene reconstruction due to the lack of temporal modeling capabilities. A common solution involves introducing time-dependent modeling through additional time conditions and deformation networks, as adopted by (Park et al. 2021a; Pumarola et al. 2020; Park et al. 2021b). Additionally, several works also improve NeRF’s reconstruction quality in both static and dynamic scenes by introducing additional supervision signals, such as semantic segmentation (Kundu et al. 2022; Zhi et al. 2021; Jain, Tancik, and Abbeel 2021) , depth maps (Xu et al. 2022; Guo et al. 2022; Deng et al. 2022a; Niemeyer et al. 2022; Deng et al. 2022b), and optical flow (Chen and Tsukada 2022; Li et al. 2023). Despite these advances, NeRF-based methods struggle with long training times and limited ability to handle complex motions. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) has demonstrated remarkable potential by representing scenes with explicit 3D Gaussian primitives, achieving both high training efficiency and superior rendering quality. Inspired by NeRF-based approaches in handling dynamic scenes, several works have extended 3DGS to model temporal changes. Specifically, (Luiten et al. 2024) proposes dynamic modeling by tracking Gaussian properties (position, orientation) while maintaining persistent appearance, while (Yang et al. 2024b) extends the 3D representation to 4D by introducing temporal Gaussian distributions. Building upon these ideas, (Yang et al. 2023b) designs dedicated deformation networks for motion modeling, (Wu et al. 2024) advances the efficiency by maintaining only canonical 3D Gaussians with HexPlane and (Bae et al. 2024) introduces embeddings to model fast and slow motions. These methods lay the groundwork for dynamic scene reconstruction in autonomous driving scenarios.
Autonomous Driving Simulation
@@
Traditional autonomous driving simulators like AirSim (Shah et al. 2018) and CARLA (Dosovitskiy et al. 2017) require extensive manual effort for environment creation while struggling to achieve photorealistic rendering. To address this, Neural based approaches have emerged as promising solutions for simulating street scenes. Early NeRF-based methods (Ost et al. 2021; Turki, Ramanan, and Satyanarayanan 2021; Rematas et al. 2022; Huang et al. 2023) introduced neural scene representation for large-scale urban environments, followed by improvements in efficiency and scalability (Tancik et al. 2022; Turki et al. 2023; Lu et al. 2023; Liu et al. 2023). For dynamic urban scene reconstruction, NSG (Ost et al. 2021) introduced supervised scene decomposition using learned scene graphs and latent object representations. This supervised approach has since become a new paradigm, adopted by MARS (Wu et al. 2023), NeuRAD (Tonderski et al. 2023) and Multi-Level Neural Scene Graphs (Fischer et al. 2024). With the advent of 3D Gaussian Splatting, DrivingGaussian (Zhou et al. 2024b) also adopted this supervised paradigm and developed hierarchical scene representations combining dynamic object graphs with incrementally updated static elements, while StreetGaussian (Yan et al. 2024) enhanced this framework with trajectory optimization and appearance modeling using 4D spherical harmonics. Recent work HUGS (Zhou et al. 2024a) further advances this line of research by incorporating physical constraints into the joint optimization of geometry, appearance, and semantics. To eliminate the need for expensive supervision, VDG (Li et al. 2024) introduced pose-free reconstruction by integrating self-supervised visual odometry for pose estimation and depth initialization, while employing motion mask supervision for static-dynamic scene decomposition. Inspired by D2NeRF’s (Wu et al. 2022) self-supervised decomposition in general dynamic scenes, SUDS (Turki et al. 2023) introduced this paradigm to autonomous driving by leveraging optical flow guidance, followed by EmerNeRF (Yang et al. 2023a) proposing a flow-free approach with hash-grid based scene organization. Despite the dominance of NeRF-based methods, this self-supervised paradigm was later extended to 3D Gaussian Splatting by several works. PVG (Chen et al. 2023) pioneered this extension by introducing periodic vibration-based temporal dynamics for unified representation of both static and dynamic elements, followed by S3Gaussian (Huang et al. 2024).
However, current methods typically employ binary static-dynamic classification, which oversimplifies the complex motion patterns in street scenes. This limitation becomes particularly evident when reconstructing objects with varying velocities, as shown in our evaluations on the Waymo-NOTR dataset. To address this challenge, we propose Explicit Motion Decomposition (EMD), a plug-and-play approach that better captures the continuous spectrum of motion in street scenes. Before presenting our method, we first review the fundamentals of 3D and 4D Gaussian Splatting.
@@
Figure 2: Overview of our Explicit Motion Decomposition (EMD) framework. Given input Gaussian primitives, our method processes them through two main components: (a) Motion-aware Feature Encoding, which combines spatial, temporal, and Gaussian-specific information to capture motion characteristics; and (b) Dual-scale Deformation Framework, which hierarchically models fast global motions and slow local deformations. The framework can be seamlessly integrated into both supervised (StreetGaussian) and self-supervised (S3Gaussian) approaches through our proposed integration strategies.
@@ 3 Preliminaries
3D Gaussian Splatting (3D-GS) (Kerbl et al. 2023) proposes an explicit rendering approach to represent 3D scenes through a collection of 3D Gaussian primitives $\mathbb{G}=\left\{\left(\mu_{k},\Sigma_{k},\alpha_{k},\mathbf{c}_{k}\right)\right\}_{k=1}^{K}$, where $K$ is the total number of Gaussians. Each Gaussian primitive represents a probabilistic distribution of density in 3D space, defined by its probability density function:
@@
where $x\in\mathbb{R}^{3}$ represents any point in the 3D world space, $\mu_{k}\in\mathbb{R}^{3}$ and $\Sigma_{k}\in\mathbb{R}^{3\times 3}$ denote the mean position and covariance matrix in world space, respectively, where the covariance matrix determines the shape and orientation of the Gaussian, $\alpha_{k}\in[0,1]$ is the opacity, and $\mathbf{c}_{k}$ encodes the view-dependent color information.
Rendering Process
For rendering, each 3D Gaussian is projected onto the image plane, where the 3D mean $\mu_{k}$ is transformed to 2D mean $\mu_{k}^{2D}$, and the world space covariance matrix $\Sigma_{k}$ is transformed to screen space as $\Sigma_{k}^{\prime}=JW\Sigma_{k}W^{T}J^{T}$, with $W$ and $J$ being the viewing transformation and projective transformation Jacobian matrices, respectively. The final pixel color at screen space position $x$ is computed through front-to-back alpha compositing:
@@
where $\alpha_{k}(x)=\alpha_{k}\exp\left(-\frac{1}{2}\left(x-\mu_{k}^{2D}\right)^{T}{\Sigma_{k}^{\prime}}^{-1}\left(x-\mu_{k}^{2D}\right)\right)$ represents the opacity contribution of the $k$-th Gaussian at pixel $x$, computed as the product of the Gaussian’s base opacity $\alpha_{k}$ and its projected 2D Gaussian evaluation. $\mathcal{N}(\mathbf{x})$ represents the set of indices of Gaussians intersecting pixel $x\in\mathbb{R}^{2}$.
In practice, for each Gaussian $k$, we parameterize its covariance matrix $\Sigma_{k}$ using rotation quaternion $\mathbf{q}_{k}$ and scaling vector $\mathbf{s}_{k}$ as:
@@
where $R(\mathbf{q}_{k})$ is the rotation matrix defined by quaternion $\mathbf{q}_{k}$, and $S(\mathbf{s}_{k})$ is the diagonal scaling matrix defined by scaling vector $\mathbf{s}_{k}$. The view-dependent color $\mathbf{c}_{k}$ is encoded using spherical harmonics (SH) coefficients:
$\mathbf{c}_{k}(\mathbf{d})=\sum_{l=0}^{L}\sum_{m=-l}^{l}k_{l,m}^{(k)}Y_{l,m}(\mathbf{d}),$ (4)
where $\mathbf{d}$ is the viewing direction, $Y_{l,m}$ are the spherical harmonic basis functions, and $k_{l,m}^{(k)}$ are the corresponding coefficients for the $k$-th Gaussian.
@@
4D Gaussian Splatting extends the static 3D-GS framework to handle dynamic scenes by incorporating temporal information. For a dynamic scene captured at different timestamps $t\in[0,T]$, each Gaussian primitive is now characterized by time-varying parameters: $\mathbb{G}(t)=\left\{\left(\mu_{k}(t),\Sigma_{k}(t),\alpha_{k}(t),\mathbf{c}_{k}(t)\right)\right\}_{k=1}^{K}$. The probability density function of each Gaussian at time $t$ becomes:
@@
The temporal evolution of these parameters is typically modeled through one of two main approaches:
Deformation-based Approach
This approach models temporal changes by applying a deformation field to the base Gaussians. For each timestamp $t$, the position of each Gaussian is updated as:
@@
Trajectory-based Approach
Alternatively, the temporal evolution can be represented by explicitly modeling the continuous trajectory of each parameter. For instance, the position trajectory can be parameterized using a set of N control points $\{\mathbf{p}_{i}\}_{i=1}^{N}$ and basis functions $\{\phi_{i}(t)\}_{i=1}^{N}$:
$\mu_{k}(t)=\sum_{i=1}^{N}\mathbf{p}_{i}\phi_{i}(t),$ (7)
where the basis functions $\phi_{i}(t)$ can be B-splines or other temporal interpolation functions. The rendering process remains similar to static 3D-GS, but now Gaussian parameters are evaluated at the specific timestamp $t$ before projection and compositing.
@@ 4 Methodology
4D Gaussian Splatting has shown promising results in dynamic scene reconstruction. However, modeling dynamic street scenes remains challenging due to diverse motion patterns. To better handle the varying motion patterns in street scenes, we propose Explicit Motion Decomposition (EMD), a plug-and-play module that can be seamlessly integrated into existing 4D Gaussian-based frameworks to enhance their capability in handling dynamic scenarios, as illustrated in Fig. 2.

### v2

@@ Abstract
Photorealistic reconstruction of street scenes is essential for developing real-world simulators in autonomous driving. While recent methods based on 3D/4D Gaussian Splatting (GS) have demonstrated promising results, they still encounter challenges in complex street scenes due to the unpredictable motion of dynamic objects. Current methods typically decompose street scenes into static and dynamic objects, learning the Gaussians in either a supervised manner (e.g., w/ 3D bounding-box) or a self-supervised manner (e.g., w/o 3D bounding-box). However, these approaches do not effectively model the motions of dynamic objects (e.g., the motion speed of pedestrians is clearly different from that of vehicles), resulting in suboptimal scene decomposition. To address this, we propose Explicit Motion Decomposition (EMD), which models the motions of dynamic objects by introducing learnable motion embeddings to the Gaussians, enhancing the decomposition in street scenes. The proposed plug-and-play EMD module compensates for the lack of motion modeling in self-supervised street Gaussian splatting methods. We also introduce tailored training strategies to extend EMD to supervised approaches. Comprehensive experiments demonstrate the effectiveness of our method, achieving state-of-the-art novel view synthesis performance in self-supervised settings. The code is available at: https://qingpowuwu.github.io/emd.
{strip}[Uncaptioned image]
Figure 1: Previous street Gaussian splatting methods find it challenging to accurately model the motion patterns of dynamic objects, which leads to blurry reconstructions. With the introduction of the proposed Explicit Motion Decomposition (EMD), which compensates for the modeling of dynamic object motion, achieving the state-of-the-art reconstruction quality.
@@ 1 Introduction
Novel view synthesis for dynamic street scenes is essential in closed-loop autonomous driving. Recent advances in neural rendering, particularly Neural Radiance Fields (NeRF) [31] and 3D Gaussian Splatting (3DGS) [19], have emerged as promising scene reconstruction methods. These pioneering approaches excel in capturing complex geometries and appearances through implicit or explicit neural representations and have been extended to dynamic scenes  [36, 39, 23, 48, 37, 56, 59, 51, 7, 3, 40, 53, 52]. Despite advancements, reconstructing autonomous driving scenes remains difficult due to complex multi-object dynamics, complex environments, and varied motion patterns.
@@
To tackle this challenge, existing works based on dynamic NeRF and 3DGS frameworks separate autonomous driving scenes into static and dynamic components through two primary paradigms: Supervised methods acquire priors for dynamic objects, such as segmentation masks [20], optical flow [45], or 3D bounding boxes from object tracking. Representative works like StreetGaussian [60] and OminiRe [6] demonstrate the effectiveness of using supervision signals for appearance modeling of dynamic objects. In contrast, self-supervised methods achieve static-dynamic separation without explicit supervision, as shown by S3Gaussian [16] and DeSiRe-GS [38], which leverage inherent motion cues to optimize a 4D street representation.
While both paradigms yield promising results, supervised methods adopt a binary classification of scene elements as either static or dynamic, which overlooks the continuous spectrum of motion inherent in street scenes. Self-supervised methods optimize the entire scene holistically but neglect the varied motion speeds among objects—for example, pedestrians generally move much slower than vehicles. Existing supervised methods [60, 6] mitigate dynamic errors by optimizing 3D bounding boxes, yet self-supervised approaches still lack an effective motion modeling mechanism.
@@
To address the different motion patterns in street scenes, we propose an Explicit Motion Decomposition (EMD) module that can be easily integrated into existing self-supervised frameworks (Fig. 2). EMD improves scene decomposition by incorporating motion-aware feature encoding and dual-scale deformation modeling. Specifically, we enhance each Gaussian primitive with learnable motion embeddings to capture its motion characteristics and design a hierarchical deformation framework that separately manages fast, global motions and slow, local deformations. This design allows for more efficient analysis of complex street scenes with varying motion speeds. We conduct extensive experiments on the Waymo and KITTI datasets by integrating EMD with representative self-supervised methods: S3Gaussian and DeSiRe-GS. In addition, EMD can be seamlessly extended to supervised methods. Our main contributions include:
@@
  • We propose EMD, the first plug-and-play module that effectively addresses varying motion speeds in street scenes through explicit motion modeling.
  • We introduce tailored training strategies for self-supervised street Gaussian splatting methods and further extend EMD to supervised settings.
  • Comprehensive experiments on Waymo and KITTI datasets demonstrate previous methods with EMD exhibit better reconstruction quality, achieving state-of-the-art (SOTA) performance in self-supervised settings.
@@ 2 Related Work
Dynamic Scene Representation. Neural Radiance Fields (NeRF) [31] revolutionizes novel view synthesis by introducing volumetric scene representation and has been improved through various extensions [1, 2, 32, 26]. However, these methods are restricted to static scenes. A common solution involves introducing additional time conditions, as adopted by [36, 39, 37]. By introducing additional supervision signals, several works [21, 66, 18, 58, 14, 9, 33, 8, 4, 24] improve the rendering quality in both static and dynamic scenes. Despite these advances, NeRF-based methods struggle with long training times and limited ability to handle complex motions. Recently, 3D Gaussian Splatting (3DGS) [19] has achieved both high efficiency and superior rendering quality by representing scenes with explicit 3D Gaussian primitives. Several works [30, 64, 62, 54] have extended 3DGS to incorporate temporal information, enabling the 4D scene representation. These methods lay the groundwork for dynamic scene reconstruction in autonomous driving scenarios.
@@
Autonomous Driving Simulation. Traditional autonomous driving simulators like AirSim [43] and CARLA [10] require extensive manual effort for environment creation while struggling to achieve photorealistic rendering. To address this approaches based on neural fields have emerged as promising solutions for simulating street scenes. Early NeRF-based methods [34, 49, 41, 17] introduce neural scene representation for large-scale urban environments, followed by improvements in efficiency and scalability [46, 50, 28, 25]. For dynamic urban scene reconstruction, several methods [34, 55, 47, 11] introduce supervised scene decomposition using learned scene graphs and latent object representations. With the advent of 3DGS, DrivingGaussian [68], StreetGaussian [60], HUGS [67] and OmniRe [6] also adopt this supervised paradigm and develop hierarchical scene representations combining dynamic object graphs with incrementally updated static elements. To eliminate the need for expensive supervision, SUDS [50] and EmerNeRF [61] leverage the optical flow as decomposition guidance. PVG [5], VDG [22], S3Gaussian [16] and DeSiRe-GS [38] extend the self-supervised setting into 3DGS by a unified representation of both static and dynamic elements.
However, existing self-supervised methods focus on the scene decomposition, overlooking the diverse motion patterns in street environments. This limitation becomes particularly pronounced when reconstructing objects moving at significantly different speeds. To tackle this challenge, we are the first to explore explicit motion modeling and introduce a plug-and-play Explicit Motion Decomposition (EMD) method. EMD enhances self-supervised approaches by capturing the continuous spectrum of motion, and it can be seamlessly extended to supervised methods.
@@
Figure 2: Overview of our Explicit Motion Decomposition (EMD) framework. Given input Gaussian primitives, our method processes them through two main components: (a) Motion-aware Feature Encoding, which combines spatial, temporal, and Gaussian-specific information to capture motion characteristics; and (b) Dual-scale Deformation Framework, which hierarchically models fast global motions and slow local deformations. The framework can be seamlessly integrated into existing supervised and self-supervised approaches.
@@ 3 Preliminaries
3D Gaussian Splatting (3DGS) [19] proposes an explicit rendering approach to represent 3D scenes through a collection of 3D Gaussian primitives $\mathbb{G}=\left\{\left(\mu_{k},\Sigma_{k},\alpha_{k},\mathbf{c}_{k}\right)\right\}_{k=1}^{K}$, where $K$ is the total number of Gaussians. Each Gaussian primitive represents a probabilistic distribution of density in 3D space, defined by its probability density function:
@@
where $x\in\mathbb{R}^{3}$ represents any point in the world space, $\mu_{k}\in\mathbb{R}^{3}$ and $\Sigma_{k}\in\mathbb{R}^{3\times 3}$ denote the mean position and covariance matrix in world space, respectively, where the covariance matrix determines the shape and orientation of the Gaussian, $\alpha_{k}\in[0,1]$ is the opacity, and $\mathbf{c}_{k}$ encodes the view-dependent color information with spherical harmonics.
For rendering, each 3D Gaussian is projected onto the image plane, where the 3D mean $\mu_{k}$ is transformed to 2D mean $\mu_{k}^{2D}$, and the world space covariance matrix $\Sigma_{k}$ is transformed to screen space as $\Sigma_{k}^{\prime}=JW\Sigma_{k}W^{T}J^{T}$, with $W$ and $J$ being the viewing transformation and projective transformation Jacobian matrices. The pixel color at screen space position $x$ is computed through alpha blending:
@@
where $\alpha_{k}(x)=\alpha_{k}\exp\left(-\frac{1}{2}\left(x-\mu_{k}^{2D}\right)^{T}{\Sigma_{k}^{\prime}}^{-1}\left(x-\mu_{k}^{2D}\right)\right)$ represents the opacity contribution of the $k$-th Gaussian at pixel $x$. $\mathcal{N}(\mathbf{x})$ represents the set of indices of Gaussians intersecting pixel $x\in\mathbb{R}^{2}$. In practice, for each Gaussian $k$, we parameterize its covariance matrix $\Sigma_{k}$ using rotation quaternion $\mathbf{q}_{k}$ and scaling vector $\mathbf{s}_{k}$ as:
@@
where $R(\mathbf{q}_{k})$ is the rotation matrix defined by $\mathbf{q}_{k}$, and $S(\mathbf{s}_{k})$ is the diagonal scaling matrix defined by $\mathbf{s}_{k}$.
@@
4D Gaussian Splatting (4DGS) [54, 63] extend the static 3DGS framework to handle dynamic scenes by incorporating temporal information. For a dynamic scene captured at different timestamps $t\in[0,T]$, each Gaussian primitive is now characterized by time-varying parameters: $\mathbb{G}(t)=\left\{\left(\mu_{k}(t),\Sigma_{k}(t),\alpha_{k}(t),\mathbf{c}_{k}(t)\right)\right\}_{k=1}^{K}$. The probability density function of each Gaussian at time $t$ becomes:
@@
Then a deformation field is applied to Gaussians. For each timestamp $t$, the position of each Gaussian is updated as:
@@ 4 Methodology
4DGS has shown promising results in dynamic scene reconstruction. However, modeling dynamic street scenes remains challenging due to diverse motion patterns. To better handle the varying motion patterns in street scenes, we propose Explicit Motion Decomposition (EMD), a plug-and-play module that can be seamlessly integrated into existing street Gaussian Splatting methods to enhance their capability in handling dynamic scenarios, as illustrated in Fig. 2.